The linear equation: y=0.15x+0.79
What is the slope (or rate of change) of this equation?

Answer by nerdybill(7384)

The linear equation: y=0.15x+0.79
.
The given equation is already in the "slope-intercept" form of:
y = mx + b
where
m is the slope
b is the y-intercept at (0,b)
.
So, slope is:
0.15
Or, if you like it as a fraction:
0.15 = 15/100 = 3/20